# Object-level Geometric Structure Preserving for Natural Image Stitching

### v2

@@ Abstract
The topic of stitching images with globally natural structures holds paramount significance, with two main goals: alignment and distortion prevention. The existing approaches exhibit the ability to align well, yet fall short in maintaining object structures. In this paper, we endeavour to safeguard the overall OBJect-level structures within images based on Global Similarity Prior (OBJ-GSP), while concurrently ensuring a good alignment performance. Our approach leverages the segment anything model [28] to extract geometric structures with semantic information. We employ triangular meshes in image transformation to protect the overall shapes of objects within images. The experimental results obtained on standard benchmarks demonstrate that OBJ-GSP outperforms existing methods in both alignment and shape preservation. Our implementation and dataset are publicly available at https://github.com/RussRobin/OBJ-GSP.
@@ 1 Introduction
Image stitching aims to align multiple images and create a composite image with a larger field of view. This method is widely utilized across diverse domains, including smartphone panoramic photography [52], robotic navigation [11], and virtual reality [1, 27]. The image stitching research community grapples with two fundamental challenges:
  • Alignment: matching corresponding points in multiple images so that they are seamlessly blended together in a panoramic image.
  • Distortion prevention: preventing structures in original images from being distorted in the panorama.
@@
In recent years, the problem of alignment has largely been addressed [40]. Methods such as APAP [58] and GSP [7] divide the images into multiple grids, compute local transformation matrices within each grid, and combine them with global transformation matrices to achieve precise alignment in overlapping regions. However, they tend to pursue the alignment of feature points excessively without recognizing the shapes of objects in the images, thus distorting the shapes and leading to unnatural stitching results, as depicted in Fig. 1 (a)-(c).
@@
Figure 1: Red boxes indicate blurriness. (a) and (b) are not aligned well. (c) GSP [7] aligns well but distorts the building. Based on this, (d) GES-GSP [14] tries to prevent distortion but still fails in this case. (f) our method protects the structure of the building by sampling on object contours extracted by SAM (e).
The main concern of image stitching is to prevent distortion on the basis of good alignment performance [14, 24, 40]. Our key insight is to globally extract the structure of objects in the input images and preserve them during stitching.
Specifically, we extract object structures by SAM [28] and preserve them by optimizing the deformation of triangle meshes. Drawing inspiration from rigid object deformations [22] in computer vision, we generate triangle meshes within each object. During image transformation, these triangle meshes tend to reach a balance between projection and similarity transformations, effectively preserving the structure of the objects. As demonstrated in Fig. 1 (f), our method excels in maintaining the overall structure of objects, preventing distortion of prominent shapes within the original images.
Our approach capitalizes on object-level geometric information, and SAM extracts object-level geometric information. As shown in Fig. 1 (e), SAM treats objects as cohesive entities, transcending the segmentation of individual lines and curves adopted in previous works [14, 24]. This allows for a more nuanced understanding of the relationships between individual geometric structures.
To summarize, the main contributions of the proposed OBJ-GSP method include:
  • We propose to preserve object shapes during image transformation to stitch a natural panorama. Our method reaches a balance between projection (to align) and similarity (to prevent distortion) transformation.
  • We introduce semantic information into image stitching through SAM [28], enabling the extraction of semantically enriched, uniformly distributed object-level image structures.
  • We generate triangular meshes within objects and constrain them as a whole to undergo similarity transformations, preserving object shapes globally. Through experiments on standard image stitching datasets, our method surpasses existing methods in both alignment and shape preservation.
@@ 2 Related work
Grid-based image stitching. Autostitch [3], a pioneering work in image stitching, matches feature points and aligns them by homography transformation. Building upon this foundation, numerous stitching algorithms partition images into grids, compute geometric transformation relationships for each grid, and combine them into a global transformation to align overlapping regions and seamlessly transit the transformation to non-overlapping areas. APAP [58], AANAP [31], and GSP [7] have evolved over time, essentially addressing most alignment problems in images. However, their grid deformation methods have no knowledge of object shapes. They pay too much attention on alignment and thus causes geometric distortion. To address this, LPC [24] proposes to leverage line-point consistency to preserve structures, introducing global collinear structures to enhance desirable characteristics for image warping. GES-GSP [14] not only preserves straight-line shapes but also protects the image’s curved structures, allowing curves to undergo similarity transformations to prevent distortion.
Nevertheless, LPC [24] and GES-GSP’s [14] extraction of straight or curved lines fails to fully capture the image’s geometric structure. Their structure preservation algorithms only protect individual lines without recognizing the overall structure. To address this, we employ semantic segmentation to extract meaningful geometric structures. We use triangular grids to protect the overall object-level geometric structure, and establish connections between dispersed geometric transformations, achieving superior results.
Geometric structure extraction. In LPC [24] and GES-GSP [14], the Line Segment Detector (LSD) [17] algorithm detects straight lines in images. GES-GSP employs edge detection methods like Canny [4] and HED [51] to identify edges, which represent the geometric structure of curves. In recent years, deep learning-based semantic segmentation [45, 43, 6, 21, 57, 8] and instance segmentation [18, 20, 2, 26] methods have advanced rapidly, capable of segmenting objects with semantic information in images. However, they require large datasets [10, 34, 47] for training and can only segment fixed categories within the dataset. Thus, large-scale pretrained encoders [5, 41, 42] have been proposed under the encoder-decoder architecture, capable of fine-tuning for downstream tasks. SAM [28], introduced in early 2023, is a large model capable of semantic segmentation for point, crop box, and text prompts. SAM’s automatic mode automatically selects point prompts, segments images, and extracts object information. We employ SAM [28] to extract object-level geometric structures and preserve these shapes during stitching. Compared to LSD [17], Canny [4], and HED [51], SAM’s extracted geometric structures carry semantic meaning, distribute evenly, and have a pleasing visual appearance. LPC [24] and GES-GSP [14] only protect independent, scattered geometric lines, while we protect objects composed of multiple geometric lines as a whole, resulting in superior performance. In the future, the accuracy and speed of SAM-type methods will both improve [53, 59], further enhancing the quality and speed of our image stitching techniques.
@@
Deep-learning based stitching and OBJ-GSP. In recent years, several methods [23, 40] like UDIS [39] have attempted to model certain image stitching steps as unsupervised deep learning problems, leading to notable advances in this field. UDIS++ [40] also addresses the distortion problem on the basis of good alignment performance, which aligns perfectly with our goals. It attempts to replace feature points extracted by traditional methods with deep learning-based backbones to reconstruct panoramas through feature map reconstruction, and they specifically design distortion-solving losses based on grid transformation. We adhere to the traditional approach in the stitching domain by preserving results in grid transformation, while UDIS++ provides a completely new deep learning-based pipeline, although currently its performance is not as good as ours. We provide a detailed comparison with UDIS++ in the supplementary material.
3 A revisit of stitching pipeline
In this section, we revisit the previous state-of-the-art stitching pipeline by GES-GSP [14]. In the next section we will show how OBJ-GSP preserves object-level structures in this pipeline.
GES-GSP uses a grid mesh to guide the image deformation, where $V$ and $E$ represent the sets of vertices and edges within the grid mesh. GES-GSP aims to find a set of deformed vertex positions, denoted as $\widetilde{V}$, that minimizes the energy function $\psi(V)$. In short, GES-GSP [14] builds upon the foundation of GSP [7] by introducing the GES term, specifically aimed at preserving the shape of lines, in addition to the Alignment term, Local similarity term, and Global similarity term within GSP [7]. In this section, we provide a brief overview of the optimization equations employed in GES-GSP [14].
@@ 3 A revisit of stitching pipeline
Alignment term extracts feature points from the input images with the goal of aligning these points after geometric transformation. For each feature point pair $(p,\Phi(p))$, $\widetilde{v}(p)$ represents the position of $p$ as a linear combination of four vertex positions, and $M$ represents the set of all feature point pairs. The algorithm linearly combines the coordinates of the four vertices of each grid to represent the position of $p$ through bi-linear interpolation. By optimizing the positions of grid vertices after geometric transformation, it aims to bring $p$ as close as possible to $\Phi(p)$. Therefore, the energy equation is defined as:
@@
GES term in GES-GSP [14] involves extracting straight lines and curve structures from the image, followed by triangulation to obtain sampled points $\widehat{V_{i}}$. These points transform to ${\widehat{V_{i}}}^{desired}$ after projection transformation and similarity transformation. To bring $\widehat{V_{i}}$ as close as possible to ${\widehat{V_{i}}}^{desired}$ after the final grid deformation, GES-GSP defines the energy term as:
$\psi_{ges}(V)=\sum_{\beta=1}^{N_{c}}\sum_{\alpha=1}^{N_{s}}{\omega_{\alpha}^{\beta}E_{\alpha}^{\beta}}.$ (4)
For a detailed description of the four energy terms mentioned above, please refer to GSP [7] and GES-GSP [14] for more information. Ultimately, the sought-after vertices are determined by minimizing the following optimization equation for two images:
$\widetilde{V}=\text{argmin}_{\widetilde{V}}(\psi_{a}(\widetilde{V})+\lambda_{l}\psi_{l}(\widetilde{V})+\psi_{g}(\widetilde{V})+\lambda_{ges}\psi_{ges}(\widetilde{V})).$ (5)
The above equation represents the optimization for two images. For a scenario involving $n$ images, such as $I_{0}$, $I_{1}$, $I_{2}$, up to $I_{n-1}$, with $I_{0}$ as the reference image and the remaining $n-1$ images transformed onto the plane of $I_{0}$, the corresponding energy terms are denoted as $\psi^{1}$, $\psi^{2}$, and so on up to $\psi^{n-1}$. The final optimization is performed as:
$\widetilde{V}=\text{argmin}_{\widetilde{V}}(\psi^{1}+\psi^{2}+\ldots+\psi^{n-1}).$ (6)
4 The proposed method
In GES-GSP [14], it is necessary to extract large-scale geometric structures, including straight lines and curves, from the entire image and preserve these structures during stitching. Therefore, GES-GSP employs an edge detection algorithm to extract edges from the image. Subsequently, the image undergoes skeletonization, noise reduction, branch pruning, and corner point removal operations. Afterward, the LSD [17] algorithm is applied to detect straight lines within the image. Post-processing techniques are then employed to handle invalid edge information, including branch pruning, re-connection of discontinuous lines, and collinearity constraints. While this conventional approach can extract some geometric structures, its results often appear meaningless to human observers and fail to capture the complete geometric information of the image.
To address this issue, our OBJ-GSP variant aims to preserve structures imbued with semantic information, making the protected structures more semantically meaningful. Moreover, the algorithm used in GES-GSP [14] tends to overemphasize certain regions, leading to an excessive weighting of these areas during the stitching process. To mitigate this, we adopt semantic segmentation methods to extract object-level geometric structures. These structures are uniformly distributed and clearly delineate the boundaries of objects in the image, enabling the preservation of meaningful geometric information in the stitched result.
Figure 2: (a) Our triangle mesh, $V_{0}$ being the center of object. (b) With every edge undergoing similarity and projection transformation, object in (a) is transformed into (b). (c) Triangle mesh with near-equilateral triangles of similar sizes across the region. (d) Triangle sampling strategy employed in GES-GSP.
After obtaining the contours, we generate a triangular mesh for each semantic object, preserving the shape of the object through similarity transformations within the triangle mesh. Unlike the As-Rigid-As-Possible (ARAP) [22] method, we simplify computational complexity by directly locating the center of the object and connecting it to sampling points on the object’s semantic boundary to form a triangular mesh. In Fig 2, $V_{0}$ represents the object’s center, while $V_{1}$ and $V_{2}$ are sampling points on the semantic boundary of the object, forming a triangular mesh with these three points. $(x_{01},y_{01})$ refer to the known coordinates of a feature point in the local coordinate plane. One vertex, $V_{2}$, of the triangle can be represented using the edges of the triangle $\vec{V_{0}V_{1}}$ and an orthogonal coordinate system obtained by rotating this edge counterclockwise by 90 degrees:
$V_{2}=V_{0}+x_{01}\overrightarrow{V_{0}V_{1}}+y_{01}\left[\begin{matrix}0&1\\ -1&0\\ \end{matrix}\right]\overrightarrow{V_{0}V_{1}}.$
@@ 4 The proposed method
$\widehat{V_{2}^{desired}}=\widehat{V_{0}}+x_{01}\overrightarrow{\widehat{V_{0}}\widehat{V_{1}}}+y_{01}\left[\begin{matrix}0&1\\ -1&0\\ \end{matrix}\right]\overrightarrow{\widehat{V_{0}}\widehat{V_{1}}}.$
@@
$E_{V_{2}}=\|\widehat{V_{2}^{desired}}-\widehat{V_{2}}\|^{2}.$
@@
$E_{\{V_{0},V_{1},V_{2}\}}=\sum_{i=0,1,2}\|\widehat{V_{i}^{desired}}-\widehat{V_{i}}\|^{2}.$
@@
$E_{{{V_{0},V_{1},V}_{2}}}=\sum_{i=1,2}\|\widehat{V_{i}^{desired}}-\widehat{V_{i}}\|^{2}.$
@@
Ultimately, our objective function is given by:
@@
For fair comparison, our parameters are identical to those of GES-GSP: $\lambda_{l}=0.75$, $\lambda_{obj}=1.5$. Our $\lambda_{obj}$ corresponds $\lambda_{ges}$ in to GES-GSP.
Table 1: We report the mean MDR to measure alignment performance, and NIQE for distortion prevention evaluation. UDIS and UDIS++ are not feature point based so we only report NIQE and leave qualitative results in supplementary material. Best results are labeled with bold text. Lower MDR and NIQE indicates better stitched panorama. Improvement row compares our proposed method and GES-GSP. Our mean improvement to GES-GSP is 3.5% in MDR and 3.8% in NIQE.
Models / Datasets Our dataset AANAP [31] APAP [58] CAVE [25]
@@
GSP [7] ECCV16 1.15296 1.06183 1.25495 0.90884
GES-GSP [14] CVPR22 1.14366 1.06213 1.24249 0.90821
OBJ-GSP (ours) 1.12229 1.05930 1.20123 0.89731
Improvement(%) 1.9 0.3 3.3 1.2
@@
UDIS [39] TIP21 3.69421 3.01517 3.69421 -
UDIS++ [40] ICCV23 3.34003 2.95493 3.56812 4.07702
GSP 2.66597 2.84241 3.4356 4.04708
GES-GSP 2.64986 2.77220 3.48713 4.03835
OBJ-GSP (ours) 2.54906 2.74965 3.39280 4.01565
Improvement(%) 3.8 0.8 2.7 0.6
Models / Datasets DFW [30] DHW [15] GES-GSP [14] LPC [24]
@@
GSP 0.98457 1.08755 1.06986 1.30562
GES-GSP 0.98034 1.05619 1.04473 1.22256
OBJ-GSP (ours) 0.97259 1.00496 0.98288 1.10622
Improvement(%) 0.8 4.9 5.9 9.5
@@
UDIS 5.74137 3.28645 5.02442 3.76994
UDIS++ 5.09680 3.23392 4.93279 3.66565
GSP 5.61905 2.75485 3.84897 4.28546
GES-GSP 5.71544 2.70838 3.79240 3.35315
OBJ-GSP (ours) 5.69104 2.60825 3.70041 3.23057
Improvement(%) 0.4 3.7 2.4 3.7
Models / Datasets REW [29] SEAGULL [33] SVA [35] SPHP [32]
Mean Distorted Residuals (MDR $\downarrow$)
GSP 1.16192 1.14467 1.51158 1.17784
GES-GSP 1.15462 1.11993 1.47197 1.14340
OBJ-GSP (ours) 1.08635 1.08296 1.47813 1.07699
Improvement(%) 5.9 3.3 -0.4 5.8
Naturalness Image Quality Evaluator (NIQE $\downarrow$)
UDIS 3.61888 4.67437 8.02090 4.35149
UDIS++ 3.67661 4.38520 7.5419 4.1248
GSP 3.18549 5.10784 7.00495 3.04781
GES-GSP 2.75234 4.69390 6.96670 2.97173
OBJ-GSP (ours) 2.81480 4.08903 6.96149 2.49712
Improvement(%) -2.2 12.9 0.5 16.0

### v4

@@ Abstract
The topic of stitching images with globally natural structures holds paramount significance, with two main goals: pixel-level alignment and distortion prevention. The existing approaches exhibit the ability to align well, yet fall short in maintaining object structures. In this paper, we endeavour to safeguard the overall OBJect-level structures within images based on Global Similarity Prior (OBJ-GSP), on the basis of good alignment performance. Our approach leverages semantic segmentation models like the family of Segment Anything Model to extract the contours of any objects in a scene. Triangular meshes are employed in image transformation to protect the overall shapes of objects within images. The balance between alignment and distortion prevention is achieved by allowing the object meshes to strike a balance between similarity and projective transformation. We also demonstrate that object-level semantic information is necessary in low-altitude aerial image stitching. Additionally, we propose StitchBench, the largest image stitching benchmark with most diverse scenarios. Extensive experimental results demonstrate that OBJ-GSP outperforms existing methods in both pixel alignment and shape preservation. Code and dataset is publicly available at https://github.com/RussRobin/OBJ-GSP.
@@ 1 Introduction
Image stitching aims to align multiple images and create a composite image with a larger field of view. This method is widely utilized across diverse domains, including smartphone panoramic photography [40], robotic navigation [8], and virtual reality [1, 19]. In recent years, the problem of alignment has largely been addressed. Methods such as APAP [45] and GSP [6] divide the images into multiple grids, compute local transformation matrices within each grid, and combine them with global transformation information to achieve precise alignment in overlapping regions. Thus, the main concern of image stitching nowadays is to prevent distortion on the basis of good alignment performance.
@@
Figure 1: Red boxes indicate blurriness. (a) and (b) are not aligned well. (c) GSP [6] aligns well but distorts the building. Based on this, (d) GES-GSP [11] tries to prevent distortion but still fails in this case. (f) our method protects the structure of the building by sampling on object contours extracted by segmentation (e).
Existing works extract lines in images are preserve them in image transformation. LPC [17] extracts and matches lines in alignment. Based on good alignment performance of GSP, GES-GSP [11] adds the similarity transformation of line structures into considerations. However, (a) they only preserves line structures, ignoring overall and object-level structures, (b) focusing only on individual lines can be quite chaotic and mislead the model (Fig. 2), (c) straight or curved structures do not exist in some scenes.
Since an important criterion for humans to judge whether an image looks natural is the naturalness of the object structures within the image, our key insight is to extract these structures and preserve them during stitching. Nowadays, state-of-the-art segmentation models can identify almost any object with superior performance. We use them to get object shapes, which represents the image structure, and then use triangle meshes to preserve these segmented object shapes during the stitching. We generate triangle meshes within each object. During image transformation, these triangle meshes tend to reach a balance between projection and similarity transformations, effectively preserving the structure of the objects. As demonstrated in Fig. 1 (f), our method excels in maintaining the overall structure of images by preventing distortion of prominent object shapes. OBJ-GSP capitalizes on object-level preserving, and we adopt leverage segmentation models to extract geometric information. As shown in Fig. 1 (e), segmentation models treats objects as cohesive entities, transcending the segmentation of individual lines and curves adopted in previous works [11, 17]. This allows for a more nuanced understanding of the relationships between individual geometric structures, and superior to previous work, it works even when there are no prominent linear structures in the images.
Previous works often used their own collected images without testing on datasets from other papers. We unified the datasets from previous works and incorporated our own collected hand-held camera and aerial images to create StitchBench, the most complete benchmark to date. We also demonstrate that in low-altitude aerial image stitching, semantic segmentation in OBJ-GSP pipeline is necessary. When the drone flies at a low altitude, the camera moves significantly, and there is a considerable distance difference between the roofs and the ground relative to the camera. These conditions do not satisfy the assumptions of image stitching [2], which assumes a fixed camera optical center or distant scenes, making stitching unfeasible. In this case, it is necessary to use a semantic segmentation model to identify the houses, then perform orthorectification to project them onto the ground before stitching.
To summarize, the main contributions of the proposed OBJ-GSP include:
  • We propose to preserve object contours before and after image transformation to maintain the overall structure of the image. Object shapes are not limited to images with obvious linear structures and are not misled by excessively noisy line structures.
  • We introduce the segmentation models into image stitching, facilitating the extraction of any object in the scene. Furthermore, we demonstrate that segmentation and OBJ-GSP are crucial for low-altitude aerial image stitching.
  • We collect StitchBench, which is by far the largest and most diverse image stitching benchmark.
@@ 2.1 Grid-based image stitching
Autostitch [2], a pioneering work in image stitching, matches feature points and aligns them by homography transformation. Building upon this foundation, numerous stitching algorithms partition images into grids, compute geometric transformation relationships for each grid, and combine them into a global transformation to align overlapping regions and seamlessly transit the transformation to non-overlapping areas. APAP [45], AANAP [23], and GSP [6] have evolved over time, essentially addressing most alignment problems in images. However, their grid deformation methods have no knowledge of object shapes. They pay too much attention on alignment and thus causes geometric distortion. To address this, LPC [17] and GES-GSP [11] propose to preserve line structures. However, (a) their method only preserves line structures, ignoring the overall structure of objects, (b) an excessive number of lines without object structure information can mislead the model, (c) some scenes do not contain straight or curved structures. We find that the large segmentation models like SAM [20] can segment all types of objects and provide their contours. This helps image stitching maintain shape consistency, so we have incorporated the family of SAM into our method. We use triangular grids to protect the overall object-level geometric structure, and establish connections between dispersed geometric transformations, achieving superior results.
2.2 Geometric structure extraction
Previous works employ Line Segment Detector [37] to detect straight lines in images, and edge detection methods like Canny [5] and HED [39] to identify edges. However, these methods require line structures to be present in the image. In cases where textures are unclear or lighting is poor, conventional methods cannot extract lines effectively, whereas large models can still operate successfully in these scenarios. We employ the family of SAM and EfficientSAM [41] to extract object-level structures and preserve them during stitching. It is notable that segmentation models are not limited by line structures and can segment almost any object. In the future, the accuracy and speed of SAM-type methods will both improve [41, 46], further enhancing the quality and speed of our image stitching techniques.
2.3 Deep-learning based stitching
@@
In recent years, several methods [18] like UDIS [30] have attempted to model certain image stitching steps as unsupervised deep learning problems, leading to notable advances in this field. UDIS++ [31] also addresses the distortion problem on the basis of good alignment performance, which aligns perfectly with our goals. We adhere to the traditional approach in the stitching domain by preserving results in grid transformation, while UDIS++ provides a completely new deep learning-based pipeline, although currently its performance is not as good as ours.
3 The proposed method
Figure 3: (a) Our triangle mesh. $V_{0}$ is the center of object. (b) With every edge undergoing similarity and projection transformation, object in (a) is transformed into (b). (c) Triangle mesh with near-equilateral triangles of similar sizes across the region. (d) Triangle sampling strategy.
OBJ-GSP introduces SAM to segment objects to obtain their structural contours and preserve object-level structures as well as aligning feature points in stitching. Locally, our approach retains the original perspective of each image. On a global scale, it seeks to preserve overall structure [6]. Moreover, at the object-level, we ensure the integrity of objects within the images, preventing distortion. To this end, we take four aspects into consideration: alignment, global similarity, local similarity and object-level shape preservation. A grid mesh is adopted to guide the image deformation, where $V$ and $E$ represent the sets of vertices and edges within the grid mesh, as shown in Fig. 3. Image stitching methods aim to find a set of deformed vertex positions, denoted as $\widetilde{V}$, that minimizes the energy function $\psi(V)$.
@@ 3 The proposed method
Alignment term extracts feature points $p$ by with an extractor (e.g. SIFT [sift]) and matches feature point pairs with matcher $\Phi$. For each feature point pair $(p,\Phi(p))$, $\widetilde{v}(p)$ represents the position of $p$ as a linear combination of four vertex positions, and $M$ represents the set of all feature point pairs. The algorithm linearly combines the coordinates of the four vertices of each grid to represent the position of $p$ through bi-linear interpolation. By optimizing the positions of grid vertices after geometric transformation, it aims to bring $p$ as close as possible to $\Phi(p)$. Therefore, the energy equation is defined as:
@@
Figure 4: APAP and SPHP see misalignment, and we delineated the indistinct portions using color-coded boxes. Autostitch, APAP, SPHP and GSP sees distortion. The convergence of the blue and red lines is essential, and we signify distortion by the intersection of these two lines. GES-GSP successfully prevents distortion, but it undergoes misalignment. Our method addresses misalignment and distortion well.
After obtaining the contours, we generate a triangular mesh for each semantic object, preserving the shape of the object through similarity transformations within the triangle mesh. Unlike the As-Rigid-As-Possible (ARAP) [15] method, we simplify computational complexity by directly locating the center of the object and connecting it to sampling points on the object’s semantic boundary to form a triangular mesh. In Fig 3, $V_{0}$ represents the object’s center, while $V_{1}$ and $V_{2}$ are sampling points on the semantic boundary of the object, forming a triangular mesh with these three points. $(x_{01},y_{01})$ refer to the known coordinates of a feature point in the local coordinate plane. One vertex, $V_{2}$, of the triangle can be represented using the edges of the triangle $\overrightarrow{V_{0}V_{1}}$ and an orthogonal coordinate system obtained by rotating this edge counterclockwise by 90 degrees:
$V_{2}=V_{0}+x_{01}\overrightarrow{V_{0}V_{1}}+y_{01}\left[\begin{matrix}0&1\\ -1&0\\ \end{matrix}\right]\overrightarrow{V_{0}V_{1}}.$ (4)
Figure 5: We magnify the ground in the red box and the car in the purple box. OBJ-GSP aligns well but the remaining five methods shows misalignment. Distortion is not observed in this case.
@@
$\widehat{V_{2}^{desired}}=\widehat{V_{0}}+x_{01}\overrightarrow{\widehat{V_{0}}\widehat{V_{1}}}+y_{01}\left[\begin{matrix}0&1\\ -1&0\\ \end{matrix}\right]\overrightarrow{\widehat{V_{0}}\widehat{V_{1}}}.$ (5)
@@
$E_{V_{2}}=\|\widehat{V_{2}^{desired}}-\widehat{V_{2}}\|^{2}.$ (6)
@@
$E_{\{V_{0},V_{1},V_{2}\}}=\sum_{i=0,1,2}\|\widehat{V_{i}^{desired}}-\widehat{V_{i}}\|^{2}.$ (7)
@@
$E_{{{V_{0},V_{1},V}_{2}}}=\sum_{i=1,2}\|\widehat{V_{i}^{desired}}-\widehat{V_{i}}\|^{2}.$ (8)
@@
To conclude, our objective function is given by:
@@
Eq. 10 can be solved with linear optimization. For fair comparison, our parameters are identical to those of GES-GSP: $\lambda_{l}=0.75$, $\lambda_{obj}=1.5$. Our $\lambda_{obj}$ corresponds $\lambda_{ges}$ in to GES-GSP.
Figure 6: OBJ-GSP for low-altitude drone image stitching. We segment our roofs and walls from aerial images first. Walls are masked out, and roofs are projected to the ground. Then OBJ-GSP stitches the projected images.
Table 1: We report the mean MDR for distortion prevention evaluation, and NIQE to measure alignment performance. UDIS and UDIS++ are not feature point based so we only report NIQE and leave qualitative results in supplementary material. Best results are labeled with bold text. Lower MDR and NIQE indicates better stitched panorama. Improvement row compares our proposed method and GES-GSP. Our mean improvement to GES-GSP is 3.5% in MDR and 3.8% in NIQE.
Models / Datasets OBJ-GSP AANAP APAP CAVE DFW DHW
@@
GSP ECCV16 1.15296 1.06183 1.25495 0.90884 0.98457 1.08755
GES-GSP CVPR22 1.14366 1.06213 1.24249 0.90821 0.98034 1.05619
OBJ-GSP (ours) 1.12229 1.05930 1.20123 0.89731 0.97259 1.00496
Improvement(%) 1.9 0.3 3.3 1.2 0.8 4.9
@@
UDIS TIP21 3.69421 3.01517 3.69421 - 5.74137 3.28645
UDIS++ ICCV23 3.34003 2.95493 3.56812 4.07702 5.09680 3.23392
GSP ECCV16 2.66597 2.84241 3.4356 4.04708 5.61905 2.75485
GES-GSP CVPR22 2.64986 2.77220 3.48713 4.03835 5.71544 2.70838
OBJ-GSP (ours) 2.54906 2.74965 3.39280 4.01565 5.69104 2.60825
Improvement(%) 3.8 0.8 2.7 0.6 0.4 3.7
Models / Datasets GES-GSP LPC REW SEAGULL SVA SPHP
@@
GSP ECCV16 1.06986 1.30562 1.16192 1.14467 1.51158 1.17784
GES-GSP CVPR22 1.15462 1.11993 1.47197 1.14340 1.04473 1.22256
OBJ-GSP (ours) 0.98288 1.10622 1.08635 1.08296 1.47813 1.07699
Improvement(%) 5.9 9.5 5.9 3.3 -0.4 5.8
@@
UDIS TIP21 5.02442 3.76994 3.61888 4.67437 8.02090 4.35149
UDIS++ ICCV23 4.93279 3.66565 3.67661 4.38520 7.5419 4.1248
GSP ECCV16 3.84897 4.28546 3.18549 5.10784 7.00495 3.04781
GES-GSP CVPR22 3.79240 3.35315 2.75234 4.69390 6.96670 2.97173
OBJ-GSP (ours) 3.70041 3.23057 2.81480 4.08903 6.96149 2.49712
Improvement(%) 2.4 3.7 -2.2 12.9 0.5 16.0